2) Solve the following system

๐‘‘๐‘ฅ

๐‘‘๐‘ก

= ๐‘ฅ + ๐‘ฆ

๐‘‘๐‘ฆ

๐‘‘๐‘ก

= โˆ’2๐‘ฅ โˆ’ ๐‘ฆย 


1
Expert's answer
2021-06-08T09:36:34-0400
"\dfrac{dx}{dt}=x+y"

"\dfrac{dy}{dt}=-2x-y"

The first equation of the system is equivalent toย 

"y=\frac{dx}{dt}-x"

Thenย 

"\frac{dy}{dt}=\frac{d^2x}{dt^2}-\frac{dx}{dt}"

Put these in the second equation of the system. We have the following differential equation:


"\frac{d^2x}{dt^2}-\frac{dx}{dt}=-2x-\frac{dx}{dt}+x"


which is equivalent to


"\frac{d^2x}{dt^2}+x=0"



Its characteristic equationย "k^2+1=0."ย 


Therefore,ย "x(t)=C_1\sin t+C_2\cos t".

Thenย 

"\frac{dx}{dt}=C_1\cos t-C_2\sin t"

ย and


"y=\frac{dx}{dt}-x"

"=C_1\cos t-C_2\sin t-C_1\sin t-C_2\cos t"

ย .Consequently, the system has the general solution:


"\begin{cases}
x(t)=C_1\sin t+C_2\cos t\\
y(t)=(-C_1-C_2)\sin t+(C_1-C_2))\cos t
\end{cases}"